OpenStudy (anonymous):

Please Help Got these questions wrong there driving me crazy ! MEDAL AND FAN !

OpenStudy (anonymous):

#1 Use the following quadratic equation: 3x2 – 8x = 48 To complete the square, you would add what number to both sides of the equation? Select one: a. 4 Incorrect b. 16 c.62/36 d. 64/36

OpenStudy (anonymous):

#2 Use the following quadratic equation: 3x2 – 8x = 48 The coordinates to the vertex would be: Select one: a. eq_3231a2.gif Incorrect b. equation c. equation d. None of the above

OpenStudy (anonymous):

#3 Use the following quadratic equation: 3x2 – 8x = 48 The value(s) of x would be: Select one: a. 2 or –2 b. 4 or –4 c. Incorrect d. 

OpenStudy (anonymous):

#4 Use the following quadratic equation: 3x2 – 8x = 48 The graph of the parabola would: Select one: a. Open downward as the middle coefficient is positive. b. Open downward as the middle coefficient is negative. Incorrect c. Open upward because the leading coefficient is negative. d. Open upward because the leading coefficient is positive.

OpenStudy (anonymous):

#5 Use the quadratic equation: x2 + 6x – 16 = 0 To complete the square, you would add what number to both sides of the equation? Select one: a. 16 b. –16 Incorrect c. 25 d. 9

OpenStudy (anonymous):

#6 Use the quadratic equation: x2 + 6x – 16 = 0 After completing the square, the quadratic equation would look like: Select one: a. (x – 3)2 = 25 b. (x – 3)2 = –25 Incorrect c. (x + 3)2 = 25 d. (x + 3)2 = 5

OpenStudy (anonymous):

#7 Use the quadratic equation: x2 + 6x – 16 = 0 The axis of symmetry would be: Select one: a. 3 b. –3 c. 6 d. No axis of symmetry as the parabola is not symmetrical. Incorrect

OpenStudy (anonymous):